OpenStudy (anonymous):

The greatest common factor of 54x and 96xy is: A. 13x B. 2x C. 9x D. 6x

OpenStudy (anonymous):

6x

OpenStudy (cwrw238):

54 = 2*3*3*3 96 = 2*2*2*2*2*3 common factors = 2*3 = 6 common factor = 6x